# Unisolver: PDE-Conditional Transformers Are Universal PDE Solvers

### v3

@@ Abstract
Deep models have recently emerged as a promising tool to solve partial differential equations (PDEs), known as neural PDE solvers. While neural solvers trained from either simulation data or physics-informed loss can solve PDEs reasonably well, they are mainly restricted to a few instances of PDEs, e.g. a certain equation with a limited set of coefficients. This limits the generalization of neural solvers to diverse PDEs, impeding them from being practical surrogate models for numerical solvers. In this paper, we present the Universal PDE Solver (Unisolver) capable of solving a wide scope of PDEs by training a novel Transformer model on diverse data and conditioned on diverse PDEs. Instead of purely scaling up data and parameters, Unisolver stems from the theoretical analysis of the PDE-solving process. Our key finding is that a PDE solution is fundamentally under the control of a series of PDE components, e.g. equation symbols, coefficients, and boundary conditions. Inspired by the mathematical structure of PDEs, we define a complete set of PDE components and flexibly embed them as domain-wise (e.g. equation symbols) and point-wise (e.g. boundaries) conditions for Transformer PDE solvers. Integrating physical insights with recent Transformer advances, Unisolver achieves consistent state-of-the-art results on three challenging large-scale benchmarks, showing impressive performance gains and favorable PDE generalizability.
@@ 1 Introduction
Partial differential equations (PDEs) are essential for numerous scientific and engineering problems (Evans, 2022, Arnol’d, 2013), such as meteorology, electromagnetism and thermodynamics (Wang et al., 2023). Since it is usually hard to obtain an analytic solution for a PDE, numerical methods are widely explored (Ames, 2014). However, these numerical methods often require huge computation costs to generate a precise solution for each PDE. Recently, deep learning models have facilitated significant advancements across a wide range of domains (Devlin et al., 2019, Liu et al., 2021, Jumper et al., 2021) and have been applied to solving PDEs, i.e. neural PDE solvers (Karniadakis et al., 2021). Owing to their excellent capability to approximate nonlinear mappings, deep models can learn to fit pre-collected data (Li et al., 2021a) or physics-informed loss function (Raissi et al., 2019) and generalize in a flash to new samples, providing an efficient approach to solving PDEs.
@@
As shown in Figure 1, previous neural solvers can be broadly categorized into two paradigms: physics-informed neural networks (PINNs) (Raissi et al., 2019) and neural operators (Li et al., 2021a). The former trains deep models using a formalized PDE loss function, while the latter solely relies on pre-collected data. However, for PINNs, while formulating the PDE equations as objective functions can ensure relatively accurate solutions, they struggle to generalize to new scenarios, necessitating retraining the model for each new task. Neural operators, on the other hand, directly learn from the data and tend to generalize better to diverse initial states and PDEs than PINNs. Nevertheless, purely based on training data may be insufficient to guide PDE solving. For example, in the case of a fluid governed by renowned Navier-Stokes equations, the typical task of neural operators is to predict future states based on past observations (Li et al., 2021a), while different viscosity coefficients and forcing terms will lead to distinct solutions even when the initial states stay the same. Thus, due to the omission of PDE information, current neural operators are mainly trained and tested on a limited set of PDEs. Notably, as neural solvers are expected to be efficient surrogate models of classical numerical solvers, generalization to various PDEs is essential for a practical neural solver.
To tackle the generalization deficiency, several works have been proposed by incorporating the PDE information into deep models or training models with a large-scale dataset. For example, message-passing neural PDE solver (Brandstetter et al., 2022) concatenates the PDE coefficients with inputs. PDEformer (Ye et al., 2024) formalizes the PDE equation as a computation graph and employs the graph Transformer (Ying et al., 2021) to aggregate PDE information. Although these methods have explored the potential of training models with both data and PDE information, they fail to consider a complete set of PDE information, thereby limiting their generalizability in some aspects. As for the other branches, such as DPOT (Hao et al., 2024), they purely scale up the training sets with diverse PDEs and expect the generalizability emerges from large data and parameters. Although models can implicitly extract PDE information from observations, the extraction process is inherently complex and resembles the challenges associated with solving inverse problems (Karniadakis et al., 2021). Therefore, these models often end up fitting an insufficient or vague representation of the underlying observation distribution, which ultimately hampers their generalizability to broader PDE solving.
@@
Figure 1: Neural PDE solvers typically consist of two paradigms: physics informed and data driven. Our proposed Unisolver combines data-driven methods with physical insights from complete PDE components in a conditional modeling framework, thereby boosting generalizability and scalability.
@@
Going beyond prior methods, as shown in Figure 1, this paper presents Unisolver as a Universal PDE solver. Concretely, Unisolver takes the advantages from both data-driven and physics-informed paradigms and empowers Transformer with favorable PDE generalizability by introducing complete physics information as conditions. Instead of simply scaling up data and parameters, we are motivated from the theoretical analysis of PDE solving and propose a complete set of PDE components. Further, drawing inspiration from the mathematical structure of PDEs, we propose to classify PDE components into domain-wise and point-wise categories according to their effect on the final solution and aggregate them as two types of deep PDE conditions. Afterward, to capture the special influence of different condition types on the hidden representations of inputs, we separate the hidden space into two subspaces and integrate these deep PDE conditions into the hidden representations of inputs in a decoupled way. We conduct extensive experiments on our own generated dataset and two large-scale benchmarks with various PDE components, where Unisolver achieves consistent state-of-the-art with sharp relative gains. Overall, our contributions are summarized as follows:
@@
  • Unisolver achieves consistent state-of-the-art performances across three large-scale benchmarks with impressive relative gains and presents favorable generalizability and scalability.
@@ 2.1 Neural PDE Solvers
Previous neural PDE solvers can be roughly categorized into the following two paradigms. The first paradigm is physics-informed neural networks (PINNs) (Raissi et al., 2019), which optimize the deep model by formalizing the PDE equations as objective functions. During training, the model outputs and gradients will gradually satisfy the targeted PDE, thereby successfully instantiating the solution as a deep model. However, it is usually hard for PINNs to generalize to unseen PDEs, limiting their applications in broader practice (Wang et al., 2023). Another booming direction is neural operators, which learn from extensive data to approximate functional dependence between input and output Banach spaces (Lu et al., 2021, Kovachki et al., 2023). Among various neural operators, FNO (Li et al., 2021a) and its variants (Li et al., 2023c, Rahman et al., 2023, Wen et al., 2022) are popular and well-established. FNO (Li et al., 2021a) effectively approximates the kernel integral operator in the frequency domain through Fourier transformation. Besides, LSM (Wu et al., 2023) generalizes the classical spectral method in latent space to tackle the curse of dimensionality. Recently, given the impressive progress achieved by Transformers (Vaswani et al., 2017), they have also been applied to solve PDEs. Existing methods treat inputs as a sequence of tokens and adopt the attention mechanism to approximate integral for solving PDEs. OFormer (Li et al., 2023a) and GNOT (Hao et al., 2023) treat each mesh point as a token and utilize the linear Transformer to avert the complexity problem. Factformer (Li et al., 2023b) axially factorizes the attention block to boost the model efficiency. Recently, Transolver (Wu et al., 2024) proposes to learn the intrinsic physical states as tokens behind input meshes, deriving a physics-attention mechanism. Despite the success of neural operators, they are only tested on the dataset containing limited PDEs. The effectiveness of these methods under large datasets containing various PDEs has not been fully explored.
@@ 2.2 Generalizable PDE Solvers
In addition to model architectures, the generalizability of neural solvers, the major advantage of numerical solvers, has also been explored. The research mainly lies in the following two directions.
@@
To guide the PDE-solving process, PDE information has been explored in many deep models. For example, PINO (Li et al., 2021b) imposes explicit equation constraints at a higher resolution to assist the learning of neural operators. CAPE (Takamoto et al., 2023) directly embeds PDE coefficients to adapt neural solvers to unseen equation coefficients. PROSE (Liu et al., 2023) and PITT (Lorsung et al., 2024) tokenize PDEs and embed mathematical expressions, enabling the transformer backbone to become aware of the underlying physics. PDEformer (Ye et al., 2024) represents the symbolic form of equations as a graph and the numeric components as nodes to optimize the processing of complex interactions between symbolic and numeric information. However, all of these methods, while incorporating equation information, do not leverage the mathematical structure of PDEs for complete and categorized embedding or integrating the prior knowledge of equation symbols within the context of natural language. In contrast, Unisolver leverages the capabilities of large language models (LLMs) (Touvron et al., 2023) to semantically embed the equation symbolic information and categorize the complete equation components based on mathematical insights, thereby facilitating the modeling of generalizable physical correlations.
@@
As a vital cornerstone of deep learning (Brown et al., 2020, He et al., 2022), recent research has also started to explore the effectiveness of large-scale training in solving PDEs. Subramanian et al. examine the scaling capabilities and transfer learning behaviors of FNO on three time-independent PDE families. MPP (McCabe et al., 2023) proposes an auto-regressive strategy to train on a broad fluid mechanics-oriented benchmark. DPOT (Hao et al., 2024) enhances MPP with a denoising method and trains a Fourier Transformer on massive PDE data comprised of 12 datasets. PDEformer (Ye et al., 2024) focuses on a 1D time-dependent PDE family and pre-trains a graph transformer on 3M samples under various equation conditions. ICON (Yang et al., 2023) trains a single neural operator capable of performing in-context learning across a wide range of differential equations. However, most of the existing methods fall short in effectively and completely integrating PDE information. This will be well addressed by Unisolver in a natural and generalizable way.
@@ 3 Unisolver
To tackle the incapability in generalization behind neural PDE solvers, we deeply dive into the PDE-solving process and present Unisolver to model the intricate interactions between initial observations and complete equation components, leading to a novel PDE-conditional Transformer model.
@@
Let $\mathcal{D}\subset\mathbb{R}^{d}$ be a bounded continuous set and =$\mathcal{M}=\{x_{1},\dots,x_{n}\}$ be an $n$-point discretization of $\mathcal{D}$, recording the coordinate information of each point. Assume that for each observation pair, we have observations of initial conditions $\mathbf{X}$ as input and target quantities $\mathbf{Y}$ as output on mesh $\mathcal{M}$ with governing PDE components $\mathcal{C}_{\text{PDE}}$ (e.g. equation symbols, coefficients, etc). The PDE-solving task is to approximate the input-PDE-output mapping $G:(\mathbf{X},\mathcal{M},\mathcal{C}_{\text{PDE}})\rightarrow\mathbf{Y}$. For instance, in fluid prediction, we need to predict the future vorticity field ($\mathbf{Y}$) from the initial field ($\mathbf{X}$), observation grid ($\mathcal{M}$), as well as PDE equations, viscosity coefficients, and force terms ($\mathcal{C}_{\text{PDE}}$).
@@ 3.1 Complete PDE Components
To enable complete modeling of the PDE, we attempt to incorporate all the underlying components that affect solutions into neural solvers. Here, we propose the complete PDE components by considering the classical vibrating string equation (2011) with fixed endpoints as a motivating example:
$\partial_{tt}u-a^{2}\partial_{xx}u=f(x,t),$ $(x,t)$ $\in(0,L)\times(0,T),$ (1a)
$u(0,t)=0,\;u(L,t)=0,$ $t$ $\in(0,T],$ (1b)
$u(x,0)=\phi(x),\;\partial_{t}u(x,0)=\psi(x),$ $x$ $\in[0,L].$ (1c)
Coefficient $a$ is determined by physical properties. $f$ represents the external force that drives the string’s vibrations. The domain geometry spans the range $[0,L]\times[0,T]$, which affects the mesh discretization in numerical solving. Eq. (1b) sets zero-displacements Dirichlet boundary conditions at the endpoints. Eq. (1c) specifies the string’s position $\phi(x)$ and velocity $\psi(x)$ at the initial. These equation components theoretically determine a unique solution. Next, we dive into the PDE-solving process of this example and explore how the complete equation components affect the final solution.
Theorem 1 (The connections between solutions and complete components in the motivating example).
$u(x,t)=\frac{1}{2}(\underbrace{\Phi(x+\text{{\color[rgb]{0.2627,0.4196,0.6510}a}}t)+\Phi(x-\text{{\color[rgb]{0.2627,0.4196,0.6510}a}}t)}_{\text{{\color[rgb]{0.6784,0.0275,0.0510}Initial position}}})+\frac{1}{2\text{{\color[rgb]{0.2627,0.4196,0.6510}a}}}\underbrace{\int^{x+\text{{\color[rgb]{0.2627,0.4196,0.6510}a}}t}_{x-\text{{\color[rgb]{0.2627,0.4196,0.6510}a}}t}\Psi(\xi)}_{\text{{\color[rgb]{0.6784,0.0275,0.0510}Initial velocity}}}\mathrm{d}\xi+\frac{1}{2\text{{\color[rgb]{0.2627,0.4196,0.6510}a}}}\underbrace{\int_{0}^{t}\mathrm{d}\tau\int_{x-\text{{\color[rgb]{0.2627,0.4196,0.6510}a}}(t-\tau)}^{x+\text{{\color[rgb]{0.2627,0.4196,0.6510}a}}(t-\tau)}}_{\text{{\color[rgb]{0.2627,0.4196,0.6510}Geometry}}}\underbrace{f(\xi,\tau)}_{\text{{\color[rgb]{0.2627,0.4196,0.6510}Force}}}\mathrm{d}\xi,$ (2)
Proof.
See Appendix D for the explicit solution and the complete proof. ∎
From the analytical solution of the above motivating example, we pinpoint that the PDE is solved under the complex interactions between a series of equation components. Specifically, the coefficient $a$ exerts a consistent influence over the domain, while the impact of the external force $f$ is imposed point-wisely. This distinction inspires us to classify the PDE components into two categories, domain-wise and point-wise, which better capture the intricate interactions.
Table 1: A detailed categorization of complete PDE components with corresponding examples.
Groups Input Domain-wise components Point-wise components
Components Initial Equation Equation Boundary External Domain Boundary
condition symbols coefficient condition type force geometry value function
Example Eq. (1c) Eq. (1a) $a$ Robin $f(x,t)$ $[0,L]\times[0,T]$ Eq. (1b)
As presented in Table 1, the complete PDE components consist of six elements. The equation formulation is defined as a domain-wise component due to its consistency across all locations. Domain geometry is categorized as a point-wise component since it is usually recorded as a binary mask and each point’s inclusion is determined individually. Boundary conditions are more complicated due to their diverse forms, e.g. periodic and Robin boundary conditions. As a result, we use two components to represent boundary conditions precisely: the boundary condition type, treated as a domain-wise component, and the boundary value function, considered as a point-wise component.

### v4

# Unisolver: PDE-Conditional Transformers Towards Universal Neural PDE Solvers

@@ Abstract
Deep models have recently emerged as promising tools to solve partial differential equations (PDEs), known as neural PDE solvers. While neural solvers trained from either simulation data or physics-informed loss can solve PDEs reasonably well, they are mainly restricted to a few instances of PDEs, e.g. a certain equation with a limited set of coefficients. This limits their generalization to diverse PDEs, preventing them from being practical surrogate models of numerical solvers. In this paper, we present Unisolver, a novel Transformer model trained on diverse data and conditioned on diverse PDEs, aiming towards a universal neural PDE solver capable of solving a wide scope of PDEs. Instead of purely scaling up data and parameters, Unisolver stems from the theoretical analysis of the PDE-solving process. Inspired by the mathematical structure of PDEs that a PDE solution is fundamentally governed by a series of PDE components such as equation symbols and boundary conditions, we define a complete set of PDE components and flexibly embed them as domain-wise and point-wise deep conditions for Transformer PDE solvers. Integrating physical insights with recent Transformer advances, Unisolver achieves consistent state-of-the-art on three challenging large-scale benchmarks, showing impressive performance and generalizability. Code is available at https://github.com/thuml/Unisolver.
Machine Learning, ICML
@@ 1 Introduction
Partial differential equations (PDEs) are essential for numerous scientific and engineering problems [8, 2], such as meteorology, electromagnetism and thermodynamics [41]. Since it is usually hard to obtain an analytic solution for a PDE, numerical methods are widely explored [1]. However, these numerical methods often require huge computation costs to generate a precise solution for each PDE. Recently, deep learning models have facilitated significant advancements across a wide range of domains [5, 24, 15] and have been applied to solving PDEs, i.e. neural PDE solvers [16]. Owing to their excellent capability to approximate nonlinear mappings, deep learning models can learn to fit pre-collected data [19] or physics-informed loss function [34] and generalize in a flash to new samples, providing an efficient approach to solving PDEs.
@@
As shown in Figure 1, previous neural solvers can be broadly categorized into two paradigms: physics-informed neural networks (PINNs) [34] and neural operators [19]. The former trains deep models using a formalized PDE loss function, while the latter solely relies on pre-collected data. However, for PINNs, while formulating the PDE equations as objective functions ensures relatively accurate solutions, they struggle to generalize to new scenarios, necessitating retraining for each new task. Neural operators, on the other hand, directly learn from data and generalize better to diverse initial states and PDEs than PINNs. Nevertheless, purely based on training data may be insufficient to guide PDE solving. For example, in the case of a fluid governed by renowned Navier-Stokes equations, the typical task of neural operators is to predict future states based on past observations [19], while different viscosity coefficients and forcing terms will lead to distinct solutions even when the initial states stay the same. Thus, due to the omission of PDE information, current neural operators are mainly trained and tested on a limited set of PDEs. Notably, as neural solvers are expected to be efficient surrogate models of classical numerical solvers, generalization to various PDEs is essential for a practical neural solver.
@@
Figure 1: Neural PDE solvers typically consist of two paradigms: physics informed and data driven. Unisolver combines data-driven methods with physical insights from PDE components in a conditional modeling framework, boosting generalizability and scalability.
To tackle the generalization deficiency, several works have been proposed by incorporating the PDE information into deep models or training models with large-scale datasets. For example, message-passing neural PDE solver [3] concatenates PDE coefficients with the inputs. PDEformer [45] formalizes the PDE as a computation graph and employs the graph Transformer [46] to aggregate PDE information. Although these methods explore the potential of training models with both data and PDE information, they do not consider a complete set of PDE information, limiting their generalizability in some aspects. As for the other branches, such as DPOT [11], they purely scale up training sets with diverse PDEs and expect the generalizability emerges from large data and parameters. Although models can implicitly extract PDE information from observations, the extraction process is inherently complex and resembles the challenges associated with solving inverse problems [16]. Therefore, these models often end up fitting an insufficient or vague representation of the underlying observation distribution, which may ultimately hamper their generalizability to broader PDE solving.
@@
Going beyond prior methods, as shown in Figure 1, this paper introduces Unisolver—a PDE-conditional Transformer progressing towards a Universal Neural PDE solver. Unisolver takes advantages from both data-driven and physics-informed paradigms and empowers Transformer with favorable generalizability by introducing complete physics information as conditions. Instead of simply scaling up data and parameters, we are motivated by the theoretical analysis of PDE solving and propose a complete set of PDE components. Drawing inspiration from the mathematical structure of PDEs, we classify PDE components into domain-wise and point-wise categories according to their effect on the final solution and aggregate them as two types of deep PDE conditions. Afterward, to capture the distinct influence of different condition types on input representations, we separate the hidden space into two subspaces and integrate these deep PDE conditions into hidden representations in a decoupled way. We conduct extensive experiments on our own generated dataset and two large-scale benchmarks with various PDE components, where Unisolver achieves consistent state-of-the-art with sharp relative gains. Overall, our contributions are summarized as follows:
@@
  • Unisolver achieves consistent state-of-the-art performances across three challenging large-scale benchmarks with impressive relative gains and presents favorable generalizability and scalability.
Figure 2: Overview of universal neural PDE solving, taking the 2D mixed PDEs in Sec 4.3 as an example. Our model is jointly trained on diverse PDEs with varied initial conditions and governing PDE components, aiming for direct generalization to unseen PDEs in downstream tasks. The “Robin boundary” in gray is a valid boundary type despite not included in the example dataset.
@@ 2.1 Neural PDE Solvers
Previous neural PDE solvers can be roughly categorized into two paradigms [43]. The first is physics-informed neural networks (PINNs) [34], which optimize deep models by formalizing PDEs as objective functions. During training, model outputs and gradients gradually satisfy the targeted PDE, thereby instantiating the solution as a deep model. However, PINNs are usually hard to generalize to unseen PDEs, limiting their broader applications [41]. Another booming direction is neural operators, which learn from extensive data to approximate functional dependence between input and output Banach spaces [27, 18]. Among various neural operators, FNO [19] and its variants [21, 33, 42] are popular and well-established. FNO effectively approximates the kernel integral operator in the frequency domain. Recently, given the impressive progress achieved by Transformers [40], they have also been applied to solve PDEs. Existing methods treat inputs as sequences of tokens and adopt the attention mechanism to approximate integrals. GNOT [10] treats each mesh point as a token and utilizes the linear Transformer to address complexity issues. FactFormer [22] axially factorizes the attention block to boost efficiency. Transolver [43] learns intrinsic physical states as tokens behind input meshes, introducing a physics-attention mechanism. Despite the success of neural operators, they are only tested on the dataset with limited PDEs. Their effectiveness for various PDEs has not been fully explored.
@@ 2.2 Generalizable PDE Solvers
In addition to model architectures, the generalizability of neural solvers has also been explored. The research mainly lies in the following two directions.
@@
To guide the PDE-solving process, PDE information has been explored in many deep models. For example, PINO [20] imposes explicit equation constraints at a higher resolution to aid neural operator learning. CAPE [38] directly embeds PDE coefficients to adapt neural solvers to unseen equation coefficients. PROSE [23] and PITT [25] tokenize PDEs and embed mathematical expressions, informing the transformer of the underlying physics. PDEformer [45] represents symbolic equations as graphs and numeric components as nodes to optimize interactions between them. However, all of these methods, while incorporating equation information, do not leverage the mathematical structure of PDEs for complete and categorized embedding or integrating the prior information of equation symbols within the context of natural language. In contrast, Unisolver leverages the capabilities of large language models (LLMs) [39] to semantically embed the equation symbolic information and categorize the complete equation components based on mathematical insights for better PDE solving.
@@
As a vital cornerstone of deep learning [4, 12], recent research has also started to explore the effectiveness of large-scale training in solving PDEs. Subramanian et al. examine the scaling capabilities and transfer learning behaviors of FNO on three time-independent PDE families. MPP [29] proposes an auto-regressive strategy to train on a broad fluid mechanics-oriented benchmark. DPOT [11] enhances MPP with a denoising method and trains a Fourier Transformer on massive PDE data comprised of 12 datasets. PDEformer [45] focuses on a 1D time-dependent PDE family and pre-trains a graph transformer on 3M samples under various equation conditions. ICON [44] trains a single neural operator capable of performing in-context learning across a wide range of differential equations. However, most existing methods fall short of effectively and completely integrating PDE information, which Unisolver will address well.
@@ 3 Unisolver
To tackle the incapability in generalization behind neural PDE solvers, we dive into the PDE-solving process and present Unisolver to model intricate interactions between initial observations and complete equation components, leading to a novel PDE-conditional Transformer model.
@@
To achieve ideal generalizability, we focus on the task of universal neural PDE solving shown in Figure 2. Let $\mathcal{D}\subset\mathbb{R}^{d}$ be a bounded continuous set and =$\mathcal{M}=\{x_{1},\dots,x_{n}\}$ an $n$-point discretization of $\mathcal{D}$ recording coordinates of each point. For each observation pair, assume we have initial condition observations $\mathbf{X}$ as input and target quantities $\mathbf{Y}$ as output on the mesh $\mathcal{M}$, with governing PDE components $\mathcal{C}_{\text{PDE}}$ (e.g. PDE symbols, coefficients) that may vary for each observation. The universal neural PDE solving task is to approximate the input-PDE-output map $G\colon(\mathbf{X},\mathcal{M},\mathcal{C}_{\text{PDE}})\rightarrow\mathbf{Y}$ across a diverse training dataset and generalize in a flash to unseen PDEs.
Figure 3: Overview of Unisolver. We universally embed all PDE components, such as domain geometries, equation symbols and boundary types into deep conditions and employ a conditional Transformer to aggregate deep conditions in the decoupled subspace.
@@ 3.1 Complete PDE Components
To enable complete modeling of the PDE, we attempt to incorporate the complete PDE components, i.e. all the underlying components that affect solutions, into neural solvers.
A motivating example
Here, we clarify the concept in the context of deep learning by considering the classical vibrating string equation with fixed endpoints as a motivating example, which can be solved explicitly, as shown in [8]. The analytical solution is provided in Appendix F.
$\partial_{tt}u-a^{2}\partial_{xx}u=f(x,t),\quad(x,t)\in(0,L)\times(0,T),$ (1a)
$u(0,t)=u(L,t)=0,\quad t\in(0,T],$ (1b)
$u(x,0)=\phi(x),\;\partial_{t}u(x,0)=\psi(x),\quad x\in[0,L].$ (1c)
From the analytical solution of the motivating example, we pinpoint that the PDE is solved through complex interactions between a series of equation components, as detailed in Table 1. These components are referred to as the complete PDE components and exhibit two key shared characteristics. Specifically, the coefficient $a$ exerts the same influence over the entire domain, while the impact of the force $f$ is imposed point-wisely. This distinction inspires us to classify these components into two categories, domain-wise and point-wise, which better capture the intricate interactions.
Table 1: Categorization of complete PDE components.
Groups Components Examples
Input Initial condition Eq. (1c)
Domain-wise components Equation symbols Eq. (1a)
Equation coefficient a
Boundary condition type Robin
Point-wise components External force $f(x,t)$
Domain geometry $[0,L]\times[0,T]$
Boundary value function Eq. (1b)
Moreover, we explain the classification of the other components shown in Table 1. The equation formulation is defined as a domain-wise component due to its consistency across all locations. Domain geometry is categorized as a point-wise component since it is usually recorded as a binary mask and each point’s inclusion is determined individually. Boundary conditions are more complicated due to their diverse forms, e.g. periodic and Robin boundary conditions. As a result, we use two components to represent boundary conditions precisely: the boundary condition type, treated as a domain-wise component, and the boundary value function, considered as a point-wise component.